Seve is training for a long distance walk. He walks for 10 km in the first week, then
each week thereafter he walks an additional 500 m. If he continues this pattern for
a year, how far does he walk:
a in the last week
b in total?

Answer by Boreal(15235)

This is the sum of an arithmetic series where a1=10, d=0.5, n=52 and n-1=51
a52=a1+(n-1)*0.5=10+0.5*51=35.5 km, in the last week
S=(n/2)(a1+a52)=26*(45.5)=1183 km, total.